OpenStudy (anonymous):

put this: (x-1)^2=12(y-1) in this form: y=a(x-h)^2+k^2

OpenStudy (saifoo.khan):

\[(x-1)^2=12(y-1)\] \[(x-1)^2=12y-12)\] \[12y = (x-1)^2 +12\] \[y = \frac{(x-1)^2 +12}{12}\] \[y = \frac{(x-1)^2}{12}+ \frac{12}{12}\]

OpenStudy (saifoo.khan):

\[y = \frac{1}{12} (x-1)^2 + 1\]